OpenStudy (anonymous):

Working mothers who earn $30,000 or less per year receive a rebate on before and after school care of 75%, plus a school bonus of $6,000 that only applies to this category of earnings. Each dollar earned above this amount and up to $40,000 reduces the rebate to 55%, and for any amount above this, the rebate drops to 30%. There are three mothers – Jade earns $40,001, Emily earns $38,000 and Lucy earns $29,550. If the cost of before and after school care is $10,000 per year, what amount of money does the person have who has the least amount of money left after paying for before and after school?

OpenStudy (jewlzme17):

x <= $30,000 => 75% rebate and $6,000 bonus $30,001 <= x <= $40,000 => 55% rebate x >= $40,001 => 30% rebate Jade earns $40,001 Emily earns $38,000 Lucy earns $29,550 Cost of before and after school care is $10,000 per year Jade = $40,001 - $10,000 + (30% of $10,000) = $30,001 + $3,000 = $33,001 Emily = $38,000 - $10,000 + (55% of $10,000) = $28,000 + $5,500 = $33,500 Lucy = $29,550 - $10,000 + (75% of $10,000) + $6,000 = $25,550 + $7,500 = $33,050 Jade, technically the highest earner, ends up with the least amount of money after paying for childcare and accounting for her allowances. She ends up with only $33,001.